A balloon is filled to a volume of 20.7L at a temperature of 27.1ºC. If the pressure in the balloon is measured to be 2.20 atm, how many moles of gas are contained inside the balloon?

Could you please solve this using the Ideal gas equation?
Answer #1

Answer:-

V = 20.7 L

P = 2.20 atm

T = 27.1 oC = 273 + 27.1 = 300.1 K

From ideal gas equation,

PV = nRT

n = PV/RT

= (2.20 x 20.7)/(0.0821 x 300.1)

= 1.85 mol